OpenStudy (anonymous):

simplify sin(x/2)cos(x/2)

OpenStudy (anonymous):

i suggest double angle formula

OpenStudy (anonymous):

recall you half angle identities.. \[\sin(\frac{ \Theta }{ 2 }) = \pm \sqrt{\frac{ 1-\cos \Theta }{ 2}}\]

OpenStudy (primeralph):